# http://dimadamen.github.io/OSNOM

### v1

@@ Abstract
As humans move around, performing their daily tasks, they are able to recall where they have positioned objects in their environment, even if these objects are currently out of sight. In this paper, we aim to mimic this spatial cognition ability. We thus formulate the task of Out of Sight, Not Out of Mind – 3D tracking active objects using observations captured through an egocentric camera. We introduce Lift, Match and Keep (LMK), a method which lifts partial 2D observations to 3D world coordinates, matches them over time using visual appearance, 3D location and interactions to form object tracks, and keeps these object tracks even when they go out-of-view of the camera – hence keeping in mind what is out of sight. We test LMK on 100 long videos from EPIC-KITCHENS. Our results demonstrate that spatial cognition is critical for correctly locating objects over short and long time scales. E.g., for one long egocentric video, we estimate the 3D location of 50 active objects. Of these, $60\%$ can be correctly positioned in 3D after 2 minutes of leaving the camera view.
Keywords:
Egocentric Video Object Tracking 3D Understanding
@@ 1 Introduction
In this paper, we make three prime [C]ontributions.
[C1] We introduce the task Out of Sight, Not Out of Mind (OSNOM) – maintaining the knowledge of where all objects are, as they are moved about and even when absent from the egocentric video stream. Egocentric views allow detailed observation of objects during interactions, e.g. the camera can look into the fridge or oven, and see exactly what was picked from the drainer in the example above. However, objects that people interact with often swiftly move out of the camera’s field of view. We focus on these challenging set of active objects that are moved by the camera wearer during the video sequence. This is distinct from the episodic memory task, e.g. [15], where the search is for one object (e.g. one’s keys), localising only when it was last seen in view. Figure LABEL:fig:teaser illustrates the OSNOM task, where the 3D location of objects and their movement is tracked throughout the video regardless of whether the objects are in view or not.
To address the OSNOM challenge, [C2] we propose an approach which lifts 2D observations to the 3D world coordinate frame – by reconstructing the scene mesh and projecting 2D detections given their depth from camera and surface estimates. We then match these lifted observations using appearance and location over time to form consistent object tracks, and keep the knowledge of objects in mind when they are out of sight. This lift, match and keep (LMK) approach allows egocentric spatio-temporal understanding which humans take for granted, yet is out of reach of current methods — knowing when an object is within reach but is out-of-view, or when in-view but occluded inside a cupboard.
[C3] We evaluate our approach on 100 long videos from the EPIC-KITCHENS dataset [6] through past and future 3D location estimations over multiple time scales. We showcase that objects are out of view for $85\%$ of frames on average. Using our LMK approach, we can correctly position $64\%$ of the objects successfully after 1 minute, $48\%$ after 5 minutes, and $37\%$ after 10 minutes. Ablations demonstrate that maintaining 3D object locations over time is critical for correctly locating moving objects, and when they are occluded or out of view.
@@ 2 Related Works
A number of tasks have been recently proposed that require 3D understanding in egocentric vision, such as jointly recognising and localising actions in a 3D map [22]. A related task to ours is Visual Query localisation in 3D (VQ3D) [15]. In VQ3D, given a query image of an object, the aim is to localise only one 3D position – when the object was last seen unoccluded and in view. Tracking is thus unnecessary (e.g. SOTA on VQ3D, EgoLoc [24], is based on retrieval and notes most objects are stationary). We include EgoLoc as a baseline showcasing its limitations for the OSNOM task.
@@
3D egocentric datasets are now becoming available [34, 15, 30, 7, 16]. Examples include Ego4D [15], which provides 3D scans and sparse camera poses for 13% of the dataset, Ego-Exo4D [16] which captures multipe first- and third-person views, and the Aria Digital Twin [30], which contains both camera poses and object segmentation masks for its two environments. EPIC-Fields [44] provides a pipeline to extract point clouds and dense camera poses from egocentric videos, and provides camera estimates for videos from the EPIC-KITCHENS dataset [6] across 45 kitchens. We use the pipeline from EPIC-Fields [44] to localise cameras in the world coordinate frame, and VISOR masks [7] to indicate active objects.
@@
Object tracking through occlusion has been investigated in 2D, where maintaining object permanence, through heuristics [17] (e.g. constant velocity [3]) or learning [39, 43], can help track assignment when occluded objects reappear. However, these works do not track out of the camera’s field of view, and their datasets targeting occlusion are short term, especially those with non-synthetic footage (e.g. TCOW [45] has a maximum video length of 464 frames).
@@
Autonomous-driving typically maintains a map of the vehicle’s surroundings [47] and tracks nearby vehicles, even when out of sight. However, whilst they maintain locations of objects through occlusion [12, 35], tracks are deleted regularly as the vehicle only has to know about objects within its vicinity.
@@
Human tracking has seen progress from 2D [53, 1, 25], to 3D [32], to 3D with motion models [20, 33, 14] which predict the location of occluded humans. Although these approaches use 3D for tracking, they usually do so in the camera coordinate frame. Some recent works have explored simultaneous reconstruction of camera motion and human pose in the 3D world coordinate frame [50, 21, 51], with [41] and [50] evaluating this concept on human tracking. [19] proposes a benchmark for tracking humans from multiple ego- and exo-centric cameras.
Our approach is related to these works for human 3D tracking. We offer the first egocentric vision work that explores tracking multiple objects in the world coordinate frame. Different from humans, objects in egocentric videos do not move by themselves, and critically move in- and out- of the camera view very frequently. We build on the human tracking approach [33], but track objects whilst taking advantage of camera localisation within the environment. Specifically, [33] tracks humans under the assumption of an external camera, thus using camera coordinates (x, y and scale) to track humans only when they are within view. Instead, our approach is tailored for egocentric videos, with significant camera motion. As objects move out of the camera’s view frequently, tracking in 3D world coordinates maintains the notion of object permanence.
@@ 3 Method - Lift, Match and Keep (LMK)
Our method operates on a single untrimmed egocentric video, $E$, recorded in an indoor environment. We aim to keep track of all objects of interest in the 3D world coordinate frame. These 3D tracks capture where all objects are all the time, even when they are not visible in the camera frame, solving the task of Out of Sight, Not Out of Mind (OSNOM).
@@
We call our method Lift, Match and Keep (LMK). We first lift 2D observations of objects to 3D, match them over time, and keep objects in mind when they are out-of-sight. Section 3.1 explains how we reconstruct the global 3D representation of the static scene along with registered camera poses and monocular depth for every frame, then use this information to lift our 2D observations into the 3D world coordinate frame. Section 3.2 then takes these lifted observations and matches them in 3D across frames using distances and visual appearance similarity. Importantly for OSNOM, we maintain 3D observed positions when the objects are out of sight. Section 3.3 details how we can use knowledge from LMK to formally define object properties with respect to the camera wearer and environment. We detail LMK next.
@@ 3.1 Lift: Lifting 2D Observations to 3D
We extract scene geometry as a 3D mesh using a classical Multi-View Stereopsis pipeline [11, 38] that runs patch matching to find dense correspondences between stereo image pairs, triangulates the correspondences to estimate depth, and fuses them together into a dense 3D point cloud with surface normals. We recover a scene mesh $\mathcal{S}$ from the dense point cloud using Poisson surface reconstruction [18]. Examples of these meshes can be seen in Figure LABEL:fig:teaser.
@@
where $K$ represents the camera’s intrinsic parameters. We denote this 3D location as $l_{n}\in\mathbb{R}^{3}$. We visualise lifting to 3D in Figure 1. Note that we represent each observation as a point in 3D following previous works [24, 15]. These 3D observations at this point are still partial and only on individual frames.
@@
3.2 Match and Keep: Matching and Keeping Lifted Observations
@@
We process the egocentric video $E$ in an online manner. While an offline approach could also be pursued, we opt to replicate the human’s spatial cognition – i.e. a person only knows of an object’s location when first encountered and this is when it is kept in mind.
Track definition. Each track $\mathcal{T}^{j}$ represents the set of observations belonging to the same object. We refer to the set of all tracks at time $t$ as $\mathcal{T}_{t}$. A track has one 3D location at each point in time, whether the object is in-sight or not, and we refer to the location of $\mathcal{T}^{j}$ at time $t$ by $L(\mathcal{T}_{t}^{j})$.
Additionally, the track has an evolving appearance representation over time. It is calculated at time $t$ using the visual appearance of the most recent $\gamma$ visual features assigned to the track. Averaging visual features increases the robustness of the representation. Restricting the average to $\gamma$ recent frames acknowledges that objects change appearance over time (e.g. a bowl may be full, then dirty, then clean over the course of the video). The appearance of the track at time $t$ is denoted $V(\mathcal{T}_{t}^{j})$.
Track initialisation. If an observation $w_{n}$ represents a new, previously unseen object, i.e. is not matched to another track using the online matching described next, we initialise a new object track with this observation. We define an initialisation function $\mathcal{I}$, which initialises a new $\mathcal{T}^{J+1}$, where $J$ tracks already exist, to the current 3D location and appearance of the observation $w_{n}$. As this is the first observation of the object, the track is projected back in time from the start of the video. $\forall t\leq f_{n}$:
@@
Track update. After a track is initialised, its visual appearance and location are updated, incorporating knowledge from new observations if they are available. We define the track update function $\mathcal{U}$. It takes a track, the observation which will be used to update the track, and the time, as input:
@@
For the earliest frame in the video which contains at least one observation, we initialise one track for each of these $\mathcal{T}_{t}=\{\mathcal{I}(w_{n})\quad\forall w_{n}\in\mathcal{W}_{t}\}$. We next iterate over every subsequent time and compare $\mathcal{W}_{t}$ to the set of trajectories at time $t-1$. Matching is based on a cost function using a combination of 3D distance and visual similarity. We follow [33] and model 3D similarity $\mathcal{\sigma}_{L}$ between an observation $w_{n}$ and a track $\mathcal{T}^{j}$ by an exponential distribution, and visual similarity $\mathcal{\sigma}_{V}$ by a Cauchy distribution:
@@
We use the Hungarian algorithm $\xi$, which computes $\Phi$ between every observation in $\mathcal{W}_{t}$ and the tracks $\mathcal{T}_{(t-1)}$ [^1] A threshold for assignment cost is set to $\alpha$. It returns a set of track assignments $A_{t}$ for time $t$, where $A_{t}^{j}=w_{n}$ indicates that the track $\mathcal{T}^{j}$ is to be assigned the observation $w_{n}\in\mathcal{W}_{t}$.
@@
We update the set of all tracks and initialise new tracks, such that:
@@ 3.3 LMK for object visibility and positioning
An object may also be referred to as Out-of-sight if it is either out-of-view or occluded (i.e. in the camera’s viewing direction but cannot be detected). If an object is occluded, it is in the camera’s coverage (i.e. ahead of the camera wearer) but cannot be detected due to being behind or inside another object.
@@
Note that the object $j$ at time $t$ may be both occluded but in-reach.
@@ 4.1 Benchmarking OSNOM
Dataset. We evaluate on long videos from the EPIC-KITCHENS [6] dataset. This dataset offers unscripted recordings of single participants. All object motions in these videos are thus a result of the camera wearer moving objects around. We utilise the publicly available 3D point clouds and dense camera poses from EPIC-Fields [44] as well as active object masks from VISOR [7] in our method and baselines. These videos are 12 minutes long on average, containing a total of 7.9M masks, which correspond to 2939 objects. We use the object semantic label only for calculating the ground truth for evaluation.
@@
For most of our results, we use masks provided by VISOR, which are interpolations of ground-truth masks. This allows us to assess LMK’s performance without accumulating errors from a detector. For completion, we also ablate these results with the usage of a semantic-free detector [40] in Section 4.4.
We randomly select 10 validation videos for hyperparmeter tuning and 100 videos for evaluation.
Benchmark task. We identify a set of frames $\mathcal{F}$ in which 3 or more objects are being interacted with. Each frame $f\in\mathcal{F}$ includes objects that are in-sight and we wish to evaluate the methods’ ability to correctly locate the 3D locations of these same objects over frames $f\pm\delta$. We compare the performance of different methods as $\delta$ increases. In total, we evaluate starting from $\mathcal{F}=3467$ frames, locations at 1M frames and 2171 objects, averaging $15k$ frames and 20 objects per video. Our benchmark will be publicly available for comparisons.
Ground truth locations. Note that there is no current egocentric dataset with 3D object annotations of dynamic objects over time. We use our 2D to 3D lifting approach presented in Section 3.1 as ground-truth locations. We assess its accuracy as follows.
@@
Figure 2: 3D Projection error. Distribution of Euclidean distance errors for the same object, at the same location, comparing $l_{n}$ to $l_{n+T}$.
@@
We select a random set of objects and the corresponding time segments when these are in the same location throughout the environment. While we do not know the ground-truth, the error between the projections from multiple views, for the same object in the same location, is another way of assessing our 3D locations. As these are multiple instances of the same object in the same location, we can measure the mean 3D error in our 3D locations. Our analysis (details in supplementary) shows that the mean 3D error is 3.5cm, with 88% of all errors smaller than 6cm and 96% of all errors smaller than 10cm (Figure 2). Given these results, we find our lifting to be sufficiently accurate to be used as ground-truth locations. This also informs our metric, where we ensure our threshold for accepting assignments is sufficiently larger than the error noted here.
@@
Evaluation metric. We define a metric called Percentage of Correct Locations (PCL), drawing inspiration from the Percentage of Correct Keypoints (PCK) [49] used to evaluate pose estimation, to evaluate the spatial alignment of objects. Traditional tracking metrics do not evaluate tracks when out of sight [2, 36, 23]. PCL instead considers a predicted 3D location correct if the Euclidean distance to its ground truth 3D location is less than a threshold $R$.

### v2

# Spatial Cognition from Egocentric Video: Out of Sight, Not Out of Mind

@@ Abstract
As humans move around, performing their daily tasks, they are able to recall where they have positioned objects in their environment, even if these objects are currently out of their sight. In this paper, we aim to mimic this spatial cognition ability. We thus formulate the task of Out of Sight, Not Out of Mind – 3D tracking active objects using observations captured through an egocentric camera. We introduce a simple but effective approach to address this challenging problem, called Lift, Match, and Keep (LMK). LMK lifts partial 2D observations to 3D world coordinates, matches them over time using visual appearance, 3D location and interactions to form object tracks, and keeps these object tracks even when they go out-of-view of the camera. We benchmark LMK on 100 long videos from EPIC-KITCHENS. Our results demonstrate that spatial cognition is critical for correctly locating objects over short and long time scales. E.g., for one long egocentric video, we estimate the 3D location of 50 active objects. After 120 seconds, 57% of the objects are correctly localised by LMK, compared to just 33% by a recent 3D method for egocentric videos and 17% by a general 2D tracking method.
@@ 1 Introduction
In this paper, we operate on egocentric videos and make three prime [C]ontributions.
[C1] We introduce the task Out of Sight, Not Out of Mind (OSNOM) – maintaining the knowledge of where all objects are, as they are moved about and even when absent from the egocentric video stream. Egocentric views allow detailed observation of objects during interactions, e.g. the camera can look into the fridge or oven, and see exactly what was picked from the drainer. However, objects often swiftly move out of the camera’s field of view. We focus on these challenging set of active objects that are moved by the camera wearer during the video sequence. Our task is to position multiple dynamic objects in 3D throughout the video, both in- and out-of-view. This is distinct from existing tasks, such as episodic memory [16], which search for the presence of an object within the video, i.e. within the camera’s field of view. Instead, the OSNOM task evaluates the locations of objects even when they are out of sight. Figure LABEL:fig:teaser illustrates the OSNOM task.
To address the OSNOM challenge, [C2] we propose a simple but effective approach that tracks objects in the world coordinate frame. Specifically, we lift observations to 3D – by reconstructing the scene mesh and projecting 2D detections given their depth from camera and surface estimates. We then match these lifted observations using appearance and location over time to form consistent object tracks, and keep the knowledge of objects in mind when they are out of sight. This lift, match and keep (LMK) approach allows spatio-temporal understanding which humans take for granted, yet is out of reach of current methods — knowing when an object is within reach but is out-of-view, or when in-view but occluded inside a cupboard.
[C3] We benchmark OSNOM on 100 long videos from the EPIC-KITCHENS dataset [7] through past and future 3D location estimations over multiple timescales. We showcase that objects are out of view for $85\%$ of frames on average. Using our LMK approach, we can correctly position $64\%$ of the objects successfully after 1 minute, $48\%$ after 5 minutes, and $37\%$ after 10 minutes, consistently outperforming recent approaches for ego [26, 61] and general [60] tracking. Ablations demonstrate that maintaining 3D object locations over time is critical for correctly locating moving objects, and when they are occluded or out of view.
@@ 2 Related Works
A number of tasks have been recently proposed that require 3D understanding in egocentric vision, such as jointly recognising and localising actions in a 3D map [24]. In Visual Query localisation in 3D (VQ3D) [16], given a query image of an object, the aim is to localise only one 3D position – when the object was last seen unoccluded and in view. Tracking is thus unnecessary (e.g. SOTA on VQ3D, EgoLoc [26], is based on retrieval and notes most objects are stationary). Recently, Zhao et al. [61] propose tracking a single object in the world-coordinate frame from RGB-D videos. They track objects in 2D, then lift these to 3D using the sensor’s depth. Another recent work [18] estimates camera pose from Dust3r [51], to predict positions of static objects like parts of the sofa. These are then tracked using rotational transforms to maintain 3D object consistency. Our approach is complementary in that we do not track objects that remain static but instead focus on tracking objects that the person has moved. Importantly, both approaches only track objects when in view. Distinctly, the OSNOM task locates objects both when in- and out-of-view, offering a complete spatial cognition of dynamic objects.
@@
3D egocentric datasets are now becoming available [37, 16, 32, 8, 17]. Examples include Ego4D [16], which provides 3D scans and sparse camera poses for 13% of the dataset, Ego-Exo4D [17] which captures multiple first- and third-person views, and the Aria Digital Twin [32], which contains both camera poses and object segmentation masks for its two environments. EPIC-Fields [48] provides a pipeline to extract point clouds and dense camera poses from egocentric videos, and provides camera estimates for the EPIC-KITCHENS dataset [7] across 45 kitchens. We use the pipeline from EPIC-Fields [48] paired with dense active object masks from VISOR [8].
Most related to 3D object tracking, [61] provides a small-scaled dataset with instance-level annotations in both 2D and 3D (4.5 hours, 250 objects in 10 environments). Instead, we use a more diverse dataset collected in an unscripted manner, with our annotations covering 25 hours, more than 2K objects in 45 environments. Similarly Ego3DT [18] is evaluated on very short videos (less than 1 minute long).
@@
Object tracking through occlusion has been investigated in 2D, where maintaining object permanence, through heuristics [19] (e.g. constant velocity [3]) or learning [42, 47], can track assignment when occluded objects reappear. However, these works do not track out of the field of view, and evaluate on short-term sequences (e.g. TCOW [49] uses sequences of maximum length of 464 frames).
@@
Autonomous-driving typically maintains a map of the vehicle’s surroundings [52] and tracks nearby vehicles, even when out of sight. However, whilst maintaining object locations through occlusion [13, 38], tracks are deleted regularly as the vehicle only has to know about its vicinity.
@@
Human tracking has seen progress from 2D [60, 1, 27], to 3D [35], to 3D with motion models [21, 36, 15] which predict the location of occluded humans. Although these approaches use 3D for tracking, they usually do so in the camera coordinate frame. Recent works have explored simultaneous reconstruction of camera motion and human pose in 3D [55, 22, 56], with [45, 55] evaluating this concept on human tracking. [20] proposes a benchmark for tracking humans from multiple ego- and exo-centric cameras.
We present the first egocentric video work to track multiple objects in the world coordinate frame. Unlike humans, objects in egocentric videos do not move autonomously and frequently enter and exit the camera view. Our approach focuses on dynamic objects, tracking them in 3D space even when out of view, preserving object permanence. We detail our approach next.
@@ 3 Method - Lift, Match and Keep (LMK)
Our method operates on a single untrimmed egocentric video, $E$, recorded in an indoor environment. We aim to keep track of all objects of interest in the 3D world coordinate frame. These 3D tracks capture the locations of objects throughout the video, even when they are not visible in the camera frame, solving the task of Out of Sight, Not Out of Mind (OSNOM).
@@
We call our method Lift, Match and Keep (LMK). We first lift 2D observations of objects to 3D (Sec 3.1), match them over time, and keep objects in mind when they are out-of-sight (Sec 3.2). We detail LMK next.
@@ 3.1 Lift: Lifting 2D Observations to 3D
We extract scene geometry as a 3D mesh using a classical Multi-View Stereopsis pipeline [12, 41] that runs patch matching to find dense correspondences between stereo image pairs, triangulates the correspondences to estimate depth, and fuses them together into a dense 3D point cloud with surface normals. We recover a scene mesh $\mathcal{S}$ from the dense point cloud using Delaunay surface reconstruction [5]. Examples of these meshes can be seen in Figure LABEL:fig:teaser.
@@
where $K$ represents the camera’s intrinsic parameters. We denote this 3D location as $l_{n}\in\mathbb{R}^{3}$. We visualise lifting to 3D in Figure 1. Note that we represent each observation as a point in 3D following previous works [26, 16]. These 3D observations are still partial and only on individual frames.
@@
3.2 Match and Keep Lifted Observations
@@
We process the egocentric video $E$ online, mimicking human spatial cognition: an object’s location is tracked only after it is first encountered and this is when it is kept in mind.
Track definition. Each track $\mathcal{T}^{j}$ represents the set of observations belonging to the same object. We refer to the set of all tracks at time $t$ as $\mathcal{T}_{t}$. A track has one 3D location at each time $t$, whether the object is in-sight or not, and we refer to this location by $L(\mathcal{T}_{t}^{j})$.
Additionally, the track has an evolving appearance representation over time. It is calculated at time $t$ using the visual appearance of the most recent $\gamma$ visual features assigned to the track. Averaging visual features enhances representation robustness. Limiting the average to $\gamma$ recent frames accounts for appearance changes over time (e.g. a bowl may be full, dirty, then clean). The track’s appearance at time $t$ is denoted as $V(\mathcal{T}_{t}^{j})$.
Track initialisation. If an observation $w_{n}$ represents a new, previously unseen object, i.e., is not matched to another track using the online matching described next, we initialise a new object track with this observation. We define an initialisation function $\mathcal{I}$, which initialises a new $\mathcal{T}^{J+1}$, where $J$ tracks already exist, to the current 3D location and appearance of the observation $w_{n}$. As this is the first observation of the object, the track is projected back in time from the start of the video. $\forall t\leq f_{n}$:
@@
Track update. Once a track is initialized, its appearance and location are updated using new observations when available. The track update function $\mathcal{U}$ takes the track, observation and time as input:
@@
Given the first frame with at least one observation, we initialise one track for each of these ${\mathcal{T}_{t}=\{\mathcal{I}(w_{n})\quad\forall w_{n}\in\mathcal{W}_{t}\}}$. We next iterate over time and compare $\mathcal{W}_{t}$ to the set of trajectories at time $t-1$. Matching is based on a cost function using a combination of 3D distance and visual similarity, as in [36]. We model 3D similarity $\mathcal{\sigma}_{L}$ between an observation $w_{n}$ and a track $\mathcal{T}^{j}$ by an exponential distribution, and visual similarity $\mathcal{\sigma}_{V}$ by a Cauchy distribution:
@@
We then use a simple Hungarian algorithm as a robust method for associations as in [59, 60, 44]. Our matching algorithm $\xi$ computes $\Phi$ between every observation in $\mathcal{W}_{t}$ and the tracks $\mathcal{T}_{(t-1)}$ [^1] A threshold for assignment cost is set to $\alpha$. It returns a set of track assignments $A_{t}$ for time $t$, where $A_{t}^{j}=w_{n}$ indicates that the track $\mathcal{T}^{j}$ is to be assigned the observation $w_{n}\in\mathcal{W}_{t}$:
@@
We update tracks and initialise new tracks, such that:
@@ 3.3 LMK for object visibility and positioning
An object may also be referred to as Out-of-sight if it is either out-of-view or occluded (i.e. in the camera’s viewing direction but cannot be detected as it is behind or inside another object).
@@
Note that an object may be both occluded but in-reach.
@@ 4.1 Benchmarking OSNOM
Dataset. We evaluate on long videos from the EPIC-KITCHENS [7] dataset. This dataset offers unscripted recordings of single participants – all object motions are thus a result of the camera wearer moving objects around. We select on 110 videos from EPIC-Kitchens, 12 minutes long on average, from 45 different kitchens, for a total of 25 hours of videos. We randomly select 10 validation videos for hyperparmeter tuning and 100 videos for evaluation. These videos contain a total of 7.9M masks, which correspond to 2939 objects. We use the object semantic label only for calculating the ground truth for evaluation.
@@
For most of our results, we use masks provided by VISOR [8], which are interpolations of ground-truth masks. This allows us to assess LMK’s performance without accumulating errors from a object detector. For completion, we also ablate these results with the usage of a semantic-free detector [43] in Section 4.4.
@@
Figure 2: 3D Projection error. Distribution of Euclidean distance errors for the same object, at one location, comparing $l_{n}$ to $l_{n+T}$.
Benchmark task. Due to the length of our videos, an exhaustive evaluation for every object from every frame is intractable. We thus select challenging key-frames $\mathcal{F}$ – these are frames with 3 or more objects being interacted with. Each frame $f\in\mathcal{F}$ includes objects that are in-sight and we wish to evaluate the methods’ ability to correctly locate the 3D locations of these same objects over frames $f\pm\delta$. We compare the performance of different methods as $\delta$ increases. In total, we evaluate starting from $\mathcal{F}=3299$ frames, locations at $603k$ frames and 2007 objects, averaging $49k$ frames and 20 objects per video. Our benchmark is publicly available for comparisons (see Project Webpage).
@@
Ground truth locations. We use our 2D to 3D lifting approach presented in Section 3.1 as ground-truth locations. We quantitatively assess the error in these locations as follows. We select a random set of objects and the corresponding time segments when these are in the same location throughout the environment. The error between the projections from multiple views, for the same object in the same location, allows assessing our 3D locations. Our analysis (details in Appendix B) shows that the mean 3D error is 3.5cm, with 88% of all errors smaller than 6cm and 96% of all errors smaller than 10cm (Figure 2). Given these results, we find our lifting to be sufficiently accurate to be used as ground-truth locations. This also informs our metric, where we ensure our threshold for accepting assignments is sufficiently larger than the error noted here.
@@
Evaluation metric. Traditional tracking metrics do not evaluate tracks when out of sight [2, 39, 25]. Thus, we define a metric called Percentage of Correct Locations (PCL), drawing inspiration from the Percentage of Correct Keypoints (PCK) [54] used to evaluate pose estimation, to evaluate the spatial alignment of objects. PCL considers a correct prediction at time $t$ if the object is correctly identified at time $t$ and its predicted 3D location is within a threshold $R$ from the ground truth 3D location. As PCL is calculated throughout time, any lost tracks are captured in the metric.